OpenStudy (anonymous):

the area of a regular octagon is 35cm^2. What is the area of a regular octagon with sides five times as long? a. 625 cm^2 b. 875 cm^2 c. 175 cm^2 d. 245 cm^2

OpenStudy (ajprincess):

Suppose the area of a regular polygon is a. Then the area of a same type of regular polygon with sides b times as long is a*b*b. Nw can u do @jfree3086?

OpenStudy (anonymous):

thank you for your help

OpenStudy (anonymous):

Find the area of a sector with a central angle f 170 degrees and a diameter of 9.1 cm. Round to the nearest tenth. Possible answers: a 122.9 cm sqrd b. 30.7 cm sqrd c. 3.4 cm sqrd

OpenStudy (anonymous):

*of 170 degrees

OpenStudy (ajprincess):

Area of a sector \[=\frac{ \theta }{ 360 }*\frac{ \pi d^2 }{ 4 }\] Here \(\theta=central~angle\) d=diameter

OpenStudy (ajprincess):

Can u find it?

OpenStudy (anonymous):

no, I still missed it.

OpenStudy (ajprincess):